Find the value of the coins after 10,20,30,and 60 years use the function you identified I'm question 3 to calculate the value
1
Expert's answer
2019-12-20T04:29:48-0500

For example, we have "x" coins and function: "f(x)=x+0.1x \cdot \,year" (that is,+10% per year).

To calculate the function for "x" coins after 10, 20, 30 and 60 years (simple interest) we will use the following formula:


"X= x + x \cdot 0.1 \cdot Y"

where Y is a period of time (10, 20, 30 and 60).

To calculate the function for "x" coins after 10, 20, 30 and 60 years (compound interest) we will use the following formula:


"X=x \cdot (1+0.1)^Y"

where Y is a period of time (10, 20, 30 and 60).